OpenStudy (anonymous):

Find the nonpermissable replacement for X in this expression; x2 dvided by 2x+6

jhonyy9 (jhonyy9):

so because the denominator is 2x+6 so than this not can be equal zero - you need to get for what value of x this 2x+6 will be equal zero

jhonyy9 (jhonyy9):

do you know why the denominator not can being equal zero ?

jhonyy9 (jhonyy9):

so because than the fraction will be undefined

jhonyy9 (jhonyy9):

can

jhonyy9 (jhonyy9):

can you make the calcule

jhonyy9 (jhonyy9):

so make 2x+6=0 than assum to both sides -6 and will get 2x+6-6=0-6 2x=-6 now divide both sides by 2 2x/2 =-6/2 x=-3 ok ?

jhonyy9 (jhonyy9):

so from this result that x not can being equal -3